### v2

# Benchmarking and Building Zero-Shot Hindi Retrieval Model with Hindi-BEIR and NLLB-E5

Abstract
Given the large number of Hindi speakers worldwide, there is a pressing need for robust and efficient information retrieval systems for Hindi. Despite ongoing research, comprehensive benchmarks for evaluating retrieval models in Hindi are lacking. To address this gap, we introduce the Hindi-BEIR benchmark, comprising 15 datasets across seven distinct tasks. We evaluate state-of-the-art multilingual retrieval models on the Hindi-BEIR benchmark, identifying task and domain-specific challenges that impact Hindi retrieval performance. Building on the insights from these results, we introduce NLLB-E5, a multilingual retrieval model that leverages a zero-shot approach to support Hindi without the need for Hindi training data. We believe our contributions, which include the release of the Hindi-BEIR benchmark and the NLLB-E5 model, will prove to be a valuable resource for researchers and promote advancements in multilingual retrieval models. The datasets from Hindi-BEIR are publicly available at Hindi-BEIR
Benchmarking and Building Zero-Shot Hindi Retrieval Model with Hindi-BEIR and NLLB-E5
Arkadeep Acharya1 thanks: Work done as an intern at IBM Research, Rudra Murthy2, Vishwajeet Kumar2, Jaydeep Sen2, 1Department of Computer Science and Engineering, Indian Institute of Technology Patna, 2IBM Research arkadeep_2101ai41@iitp.ac.in,{rmurthyv,vishk024,jaydesen}@in.ibm.com
1 Introduction
Information retrieval (IR) models are indispensable in our digital age, enabling swift and accurate extraction of relevant data from vast amounts of information. These models enable quick and accurate retrieval of relevant data, thereby facilitating informed decision-making across various downstream applications and domains. In this modern era of LLMs, retrievers have become even more relevant and almost necessary in curbing hallucinations from Large Language Model (LLM) generations by means of much popular retrieval augmented generation (RAG) Lewis et al. (2021) methods. However, the focus of information retrieval (IR) research has predominantly centered on the English language, which can be attributed to the availability of comprehensive benchmarks, such as the BEIR Thakur et al. (2021). Although, more recently, there have been efforts to build multi-lingual retrievers and associated evaluation benchmarks  Zhang et al. (2023, 2021) on non-English languages, multi-lingual expansion of robust benchmarks is still a work in progress, with many languages like Hindi lacking a BEIR like a comprehensive benchmark.
Hindi is one of the official languages of India [^1] https://en.wikipedia.org/wiki/Languages_of_India, a language of immense importance and global reach, with over half a billion speakers worldwide and ranking as the third most widely spoken language [^2] https://www.thecollector.com/what-are-the-most-spoken-languages-in-the-world/. However, from NLP research point of view, Hindi can be categorized as a low-resource language, lacking comprehensive resource and benchmarks to advance scientific research and this includes lack of a good retrieval benchmark too. Thus far, the evolution of Hindi or, in general, Indic language retrieval datasets has been rather ad-hoc, restricted to a small subset of domains, languages and tasks. Although a recent effort attempts to incorporate multiple Indic languages Haq et al. (2023) for a retrieval benchmark, the absence of diversity in domains and tasks remains a major limitation, affecting robust evaluation of the current state of retriever models in practical applications, which often involve varying tasks and domains.
In this work, we posit that creating a BEIR-like diversified and robust benchmark is the necessary first step to track and advance any retrieval research effectively in a new language. Such a benchmark will be more useful than ad-hoc datasets with multiple languages. Thus, we choose the most widely spoken language among Indic languages, Hindi, as our pivot language to pioneer a large-scale diversified retrieval benchmark Hindi-BEIR, spanning 15 diverse datasets across $7$ tasks. Hindi-BEIR aims for two key objectives:(1) Establish a standardized retrieval benchmark to assess, compare, and advance the state-of-the-art retriever models and (2) Provide workable insights into the potential research directions on building retrieval models for Hindi. With multiple tasks and domains in Hindi-BEIR, we posit that Hindi-BEIR will provide a more realistic and accurate assessment of retriever models for zero-shot usage in unseen domains and applications.
On the choice of Hindi language it is important to note here, besides being a language of high usage, a retrieval benchmark in Hindi offers some unique challenges too such as:
$\bullet$ Script Difference: Hindi uses the Devanagari script, which is fundamentally different from the Latin script used in English. This affects character encoding, text normalization, and processing. Existing tokenizers trained on English data may not handle Hindi text well, thus needing to evaluate various tokenization strategies too.
[…]
$\bullet$ Ambiguity: In Hindi, some proper nouns can also function as common nouns. For instance, the name Lata, a common female name, can also refer to creeper, a common noun. A query like lata ko kaise saaph karen (how to clean a creeper) can easily mislead lexicon-based retrieval systems and test a model’s ability for word sense disambiguation.
The introduction of Hindi-BEIR benchmark lets us assess and compare the existing multi-lingual retriever models more accurately for their zero-shot performance across diverse domains and tasks in Hindi-BEIR. Based on this performance comparison, we notice that a key limitation of existing models is the need for a substantial amount of language-specific training data, which is often not available for low-resource languages such as Hindi. To address this problem, we propose NLLB-E5, a multilingual retrieval model, inspired by recent works like Langbridge Yoon et al. (2024) and NLLB-LLM2Vec Schmidt et al. (2024), which distills information from a monolingual retrieval model into a multilingual retrieval model in a zero shot setup via a multi-lingual encoder. The key strength of NLLB-E5 is it does not need multilingual training data for distillation and thus eliminates the need for supervised multilingual training data.
We summarize our contributions as follows:
  • We introduce Hindi-BEIR, the first comprehensive retrieval benchmark in the Hindi language, encompassing 15 diverse datasets from 7 distinct tasks. This was accomplished through the harmonization of high-quality quality verified translated data from BEIR, the conversion of existing Hindi datasets for retrieval purposes, and the introduction of strategically generated synthetic data.
  • We develop NLLB-E5, which combines a multilingual encoder with a monolingual retrieval model through knowledge distillation to develop a multilingual retrieval model without the need for any multilingual training data.
  • We evaluate NLLB-E5 on Hindi-BEIR and compare against existing multilingual models, which have been trained on multilingual data. We empirically demonstrate that NLLB-E5 achieves strong performance on Hindi-BEIR, outperforming existing multilingual models without the need for multilingual training data.
2 Related Works
One of the known sources of testing retriever performance on Indic languages was to test on the Indic language subset contained in multi-lingual IR datasets such as MIRACL Zhang et al. (2022), MR.Tydi Zhang et al. (2021) having instances from Indic languages such as Hindi, Bengali, and Telugu. However, two major drawbacks here were: (1) all of these subsets were specific to a domain, hence couldn’t really be used as a robust benchmark, and (2) the Indic language-specific subsets of instances didn’t contain enough data points to facilitate building neural retrieval models for Indic languages either.
More recently, a relatively large-scale retrieval benchmark MSMarco Nguyen et al. (2016) got translated into multiple languages to produce mMARCO Bonifacio et al. (2022), including Hindi from the set of Indic languages. mMARCO was further extended to more Indic languages by IndicIRSuite Haq et al. (2023), which is the most recent retrieval benchmark in Indic languages. However, the common problem for both mMARCO and IndicMarco is that these benchmarks, too, are tied to a single type of retrieval task and web domain as was originally in MSMarco and, therefore, do not provide a robust benchmark.
In terms of Language specific benchmarks, Snegirev et al. (2024) introduce the ruMTEB Benchmark, an embedding model evaluation benchmark in Russian with Retrieval being a subtask, Valentini et al. (2024) introduce the Spanish IR dataset while Wojtasik et al. (2024) release the IR Benchmark in the Polish Language. In Indian languages such as Hindi, the Forum for Information Retrieval Evaluation (FIRE) [^3] https://fire.irsi.org.in/fire/2024/home has released numerous sub-tasks and datasets over the years to facilitate the benchmarking and development of information retrieval models. However, these datasets have predominantly been sourced from various news outlets and are not freely or openly accessible to the public.
To the best of our knowledge, Hindi-BEIR presents the first comprehensive IR benchmark, which spans diverse domains and tasks and therefore, provides the first BEIR Thakur et al. (2021) equivalent comprehensive retrieval benchmark in Hindi.
For multilingual retrieval model development, the most promising works are from Chen et al. (2024) and Wang et al. (2024c), but both these models are heavily dependent on multilingual data and do not scale for languages and domains where data is not readily available in the target language.
Our NLLB-E5 model is inspired by the works by Yoon et al. (2024) who propose Langbridge, an approach which enables multilingual reasoning without multilingual supervised training and Schmidt et al. (2024) who opt for a distillation based approach and introduce NLLB-LLM2Vec capable of producing robust multilingual embeddings extending from NLLB encoder.
3 Hindi-BEIR Retrieval Benchmark
Dataset Name Tasks #Corpus # Queries Average # words
Corpus Query
ArguAna Argument Retrieval 7763 1194 159.20 178.20
FiQA-2018 Question-Answering 48178 5924 118.02 15.23
TREC-COVID Bio-Medical IR 76492 49 159.77 14.39
SCIDOCS Citation-Prediction 22050 850 169.88 12.82
SciFact Fact-Checking 2849 1099 164.37 19.44
Touché-2020 Argument Retrieval 355273 49 351.47 8.58
NQ Question Answering 2595865 2952 89.83 9.63
FEVER Fact Checking 5362876 120075 88.86 9.48
Climate-FEVER Fact Checking 5362911 1499 88.86 24.40
CC News Retrieval News Article Retrieval 5005483 49699 272.20 9.30
Sangraha-IR Question Answering 350000 9744 308.13 30.91
MIRACL Question Answering 3494 350 66.49 10.42
IndicQARetrieval Question Answering 261 1544 480.66 10.88
mMARCO Passage Retrieval 8841823 6980 64.40 6.46
WikiPediaRetreival Question Answering 13500 1500 77.19 9.54
Table 1: Statistics of the Dataset in the Hindi-BEIR Benchmark showing the tasks which each dataset covers and the number of corpus and query in the evaluation set of each dataset in the Hindi-BEIR Benchmark.
Table 1 summarises the various datasets included in the Hindi-BEIR benchmark, along with the number of documents, queries, and domain information.
To ensure that the Hindi-BEIR Benchmark is comprehensive, challenging, and accessible to the public for future research and evaluation, we adhered to the following objectives:1) Diverse Domains and Tasks: We include datasets from diverse domains like Wikipedia and news articles to niche domains like scientific publications, finance to test the robustness and generalization ability of the retrieval models. Additionally, a good retrieval model should handle documents and queries of varying lengths equally well. We have ensured that the datasets included in the Hindi-BEIR benchmark exhibit a wide range of query and document lengths among the datasets. 2) Difficulty Level: We ensure that the datasets are challenging and systems relying solely on lexical overlap have a hard time retrieving the correct document. 3) Public Availability: All datasets curated in Hindi-BEIR Benchmark have friendly, user-friendly licenses and will be made publicly available.
We discuss the Hindi-BEIR benchmark creation in the following subsections. Please refer to Appendix A.2 for a more detailed analysis of each dataset.
3.1 Translating English Datasets to Hindi:
We translate a subset of the existing English datasets from the BEIR benchmark Thakur et al. (2021) into Hindi. This approach was necessitated by the need to ensure comprehensive coverage across multiple domains and to maintain a high level of data quality and complexity. We utilized the Indic-Trans2 model  [^4] Our choice for IndicTrans2 over other translation models has been discussed in A.3 Gala et al. (2023), a multilingual NMT model supporting translations across all 22 scheduled Indic languages (including English). We employ the back-translation technique to retain good translations. Specifically, given an English query/document we translate it to Hindi. This Hindi-translated query/document is then translated back to English. We calculate the Chrf(++) score Popović (2017) between the original English query/document and the backtranslated English query/document. We retained only those translations with a Chrf(++) score exceeding a threshold. We empirically set the threshold to 50 after manually verifying the translation quality of texts obtained from different thresholds.
This strategy enables us to leverage the wealth of existing high-quality datasets in English while making them accessible and useful for Hindi language information retrieval tasks. 9 out of the 15 datasets, which include Arguana, FiQA-2018, TREC-COVID, SCIDOCS, SciFact, Touché-2020, NQ, FEVER, and Climate-FEVER were created by this method.
Figure 1: A comprehensive overview of our training methodology. The portion above the dotted lines represents the teacher model (enclosed by a blue dotted box), while the portion below represents our proposed NLLB-E5 student model (enclosed by a red dotted box). We use NLLB as a multilingual encoder and train the Linear projection layer and the Lora adapters by forcing its outputs to match that of the teacher model via the Mean Squared Error (MSE).
3.2 Using existing datasets for Retrieval task
We create Hindi CC News Retrieval and Sangraha-IR datasets from existing datasets. We detail the process we followed for each of the datasets as follows:
Hindi CC News Retrieval : This is a cross-lingual dataset derived from the Hindi subset of the Multilingual CC News dataset [^5] https://huggingface.co/datasets/intfloat/multilingual_cc_news, which contains 7,444,584 data points comprising Hindi news articles and their corresponding titles. The article text, which is in Hindi, serves as the corpus to be retrieved. At the same time, the English news titles act as queries, representing a practical and common real-world scenario that needs the models to have language-agnostic knowledge across both Hindi and English.
We selected only those title-article pairs where the title never appears in the article text, thus reducing the dataset size to 451,803. This ensures no lexical overlap between the title and the article, making the retrieval task challenging as the model needs to rely on cross-lingual cues. From these filtered data points, we randomly selected 49,699 queries to form the test split, which has been incorporated into the Hindi-BEIR benchmark. Additionally, we release training and validation splits consisting of 301,578 and 100,526 query-corpus pairs, respectively. The train, test, and validation splits collectively encompass all data points obtained after the second filtering step.
Sangraha-IR :This dataset builds on top of the Sangraha Dataset Khan et al. (2024). We handpicked 350,000 entries from the Sangraha-verified subset, focusing on topics that truly represent Indian culture, such as agriculture, Bollywood, cricket, and other similar areas that capture the heart of India. We specifically chose the Sangraha-verified corpus as it consists of scraped data from "human-verified" Websites, OCR-extracted data from high-quality Indic language PDFs, transcribed data from various Indic language videos, podcasts, movies, courses, etc, and truly captures the essence of the language. To construct a set of relevant queries, we randomly selected 10,000 data points and used the Google Gemini-1.5-flash model to generate corresponding queries  [^6] Please refer to the Appendix for the prompt used for these documents. Queries that did not adhere to the provided instructions or deviated from the desired format were excluded, resulting in a final dataset of 9,744 queries and 350K candidate documents.
3.3 Subset from existing Multilingual Datasets
We also include 5 publicly available datasets, namely MIRACL (dev split)Zhang et al. (2021), MLDR Chen et al. (2024), mMarco Bonifacio et al. (2022), IndicQARetrieval (developed by modifying the IndicQA Dataset Doddapaneni et al. (2023)) and WikiPediaRetrieval  [^7] https://huggingface.co/datasets/ellamind/wikipedia-2023-11-retrieval-multilingual-queries.
4 NLLB-E5: multilingual retriever
In this section, we build a strong multilingual retrieval model based on the approach proposed by Yoon et al. (2024) and Schmidt et al. (2024).
4.1 LangBridge
Yoon et al. (2024) proposed an approach to equip existing LLMs with multilingual capability without any multilingual supervised data. They vertically stack a multilingual encoder and an LLM, adding a projection/alignment layer between them. The input first passes through the multilingual encoder, the representation from the multilingual encoder passes through the alignment layer and then the output from the alignment layer is passed as input to the LLM. They use English-labeled data to fine-tune the overall model and update only the embedding layers and the alignment layer. The model showed better multilingual capabilities.
4.2 NLLB-LLM2vec
Schmidt et al. (2024) extended the LLM2vec model BehnamGhader et al. (2024) by vertically stacking a NLLB model Team et al. (2022) and LLM2vec model. Similar to LangBridge Yoon et al. (2024), they add a projection/alignment layer between the two and add LoRA parameters on the LLM. Different from LangBridge, they have a teacher model and a student model. They use LLM2vec as the teacher model and the NLLB-LLM2vec as the student model. The input is passed through both the teacher and the student models. They minimize the mean squared error between the final representations obtained from both the teacher and the student model. They train the model in two stages, the first stage being general alignment between the teacher and the student model, and, the second stage being task alignment. They show that the NLLB-LLM2vec model is able to obtain superior performance on various natural language understanding tasks.
4.3 NLLB-E5
Figure 1 shows the architecture of NLLB-E5 with the different components and data flow between them. We build NLLB-E5 on the back of two components (1) A multilingual encoder NLLB and (2) a monolingual retriever E5. We bank on the multilingual encoder to project semantically similar sentences across languages to a shared representation space. We hypothesize that embeddings coming from such a language-agnostic shared representation space can act as a strong prior that can be easily aligned for producing multilingual task-specific embeddings. Moreover, given the multilingual nature of the embedding coming directly from the multilingual encoder, the aforementioned task-specific alignment is language agnostic and, therefore, can be done with task-specific data in a high-resource language, too.
Figure 1 demonstrates how we follow the above intuition in designing NLLB-E5. We use NLLB Team et al. (2022) as the strong multilingual encoder that produces robust multilingual embeddings owing to its translation capabilities for over 200 languages. To align these embeddings for the retrieval task, we learn a projection layer from the multilingual NLLB encoder to a monolingual retriever E5. The alignment happens by learning the projection layer parameters between NLLB and E5 through distillation from E5 embeddings, which are specifically tuned for retrieval tasks. Since this alignment to the retrieval task is agnostic to a specific language, we use English data and monolingual retriever E5 to do the distillation, eliminating the need for Hindi training data. We chose E5-large as the teacher and student model coupled with NLLB owing to its satisfactory retrieval performance on the BEIR benchmark.
5 Training Methodology
The NLLB-E5 model is constructed by integrating the monolingual E5 model Wang et al. (2024a) atop the multilingual encoder from NLLB Team et al. (2022), connected via a learnable projection layer. The parameters of both the monolingual and multilingual base models remain frozen throughout training. To enable adaptation, we introduce $W\in{R}^{n\times d}$, a linear projection layer mapping the $n$-dimensional output space of the multilingual encoder to the $d$-dimensional input space expected by the E5 model.
We employ knowledge distillation Hinton et al. (2015) to transfer the capabilities of the monolingual E5 model to the NLLB-E5 model, facilitating multilingual sentence representation learning. Specifically, the original E5 model serves as the teacher, producing high-quality English embeddings, while the NLLB-E5 model, which includes the learnable Linear projection $W$, acts as the student. By minimizing the discrepancy between the teacher and student embeddings, the student model learns to approximate the E5 model’s performance across Hindi and other languages supported by NLLB. The frozen multilingual encoder ensures effective cross-lingual transfer, enabling generalization across all NLLB supported languages.  [^8] We also experiment with LoRA adapters added on top of the E5 model. Details have been discussed in the Appendix A.5
5.1 Training Process
The training procedure consists of the following steps:
Get the Teacher Model Embeddings: Given an input English sentence $S$, we first tokenized it using the E5 tokenizer, yielding a sequence of tokens $\{t_{1},t_{2},\dots,t_{n}\}$. These tokens are passed through the teacher E5 model, with frozen parameters, denoted as $\mathcal{M}_{T}$. The model produces a sequence of token-level representations for the tokens:
$\{X_{1},X_{2},\dots,X_{n}\}=\mathcal{M}_{T}(\{t_{1},t_{2},\dots,t_{n}\})$
where $X_{i}\in{R}^{d}$ represents the $d$-dimensional embedding of the token $t_{i}$ produced by the model $\mathcal{M}_{T}$. To obtain the final sentence-level embedding $\vec{X}_{T}$, we apply average pooling over the token-level representations, which act as the final embedding for the sentence $S$ obtained from the teacher model.
Get Student Model Embeddings: The input sentence is first tokenized using the NLLB tokenizer, producing a sequence of tokens $\{t_{1},t_{2},\dots,t_{m}\}$. The frozen multilingual encoder processes this token sequence, generating token-level representations in the multilingual semantic space $\{H_{1},H_{2},\dots,H_{m}\}$. To align with the input structure expected by the E5 model, the token sequence is modified by prepending a token embedding corresponding to either [CLS] query: (for queries) or [CLS] passage: (for corpus), obtained from the E5 embedding layer ($H_{\emph{prior}}$), along with the representation of [SEP]($H_{\emph{post}}$) which is appended at the end. This is projected through a learnable layer $W$ as $\{H^{\prime}_{1},H^{\prime}_{2},\dots,H^{\prime}_{m}\}=W\{H_{\emph{prior}},H_{1},H_{2},\dots,H_{m},H_{\emph{post}}\}$ This projected sequence is then passed through the student E5 model, which outputs the token-level representations $\{X^{\prime}_{1},X^{\prime}_{2},\dots,X^{\prime}_{m}\}$ Finally, average pooling is applied over these token-level representations to obtain the final sentence-level embedding $\vec{X}_{S}$ from the student model.
Minimizing the Loss Function. To align the student model with the teacher, we minimize the Mean Squared Error (MSE) between the teacher and student embeddings, $\vec{X}_{T}$ and $\vec{X}_{S}$, respectively, for a batch of sentences.
This optimization encourages the student to generate multilingual embeddings that are aligned with the high-quality English embeddings produced by the teacher.
6 Experiments
This section provides a detailed overview of our experiments on Hindi-BIER datasets, including our experimental setup and the baseline models used in our experimentation.
6.1 Benchmark and Evaluation Metric
We use Hindi-BEIR as the target benchmark for all our experiments. We evaluate the models across all $15$ datasets in Hindi-BEIR and compare their performance at the dataset level too for finer insights.
Following BEIR Thakur et al. (2021) benchmark standard, we use Normalised Discounted Cumulative Gain (NDCG) Järvelin and Kekäläinen (2002), more specifically NDCG@10 as our evaluation metric. NDCG is known to be a more robust metric than simple recall in the IR community because it is purposefully designed to be sensitive towards ranking of the retrieved results.
6.2 Implementation Details of NLLB-E5
To make NLLB-E5 produce robust generalizable embeddings across different kind of retrieval tasks and domains, we rely on using diverse data to learn NLLB and E5 alignment during training. We followed a similar approach as taken by generic text-to-embedding models and used a subset of data from the huge pool of  1B diverse data curated for training sentence transformer models SIMOULIN (2021). We randomly selected 100,000 samples from each dataset to create the final training set. The datasets chosen are i) Arxiv ii) Natural Questions (NQ) iii) HotpotQA iv) Stack Exchange - Duplicate questions (titles) v) Stack Exchange - (Title, Body) pairs vi) Stack Exchange - (Title+Body, Answer) pairs vii) Stack Exchange - (Title, Answer) pairs viii) Stack Exchange - Duplicate questions (bodies) ix) Stack Exchange duplicate questions (titles+bodies) x) SQuAD 2.0 xi) S2ORC Citation - pairs (Abstracts) xi) S2ORC (Title, Abstract) xii) SearchQA xiii) PubMed (Title, Abstract) xiv) SPECTER xv) FEVER xvi) Wikipedia Sections xvii) Stack Exchange Math xviii) Stack Overflow Posts xix) Wikipedia
We train the NLLB-E5 model for 10 epochs, with a learning rate of 2e-4 and a linear scheduler for learning rate adjustment with FP16 precision. We chose the best checkpoint based on the evaluation result on the MsMARCO dev set. We ran all our experiments using Nvidia A100 80GB GPUs. Evaluating the NLLB-E5 model using 4 A100 80GB GPUs on the Hindi-BEIR benchmark took 24 hours.
Dataset Name
BGE-M3
mE5-Base
mE5-Large
LASER
LaBSE
BM-25 NLLB-E5 (3.3B)
ArguAna 53.81 49.96 54.77 11.27 32.90 43.75 57.68
FiQA-2018 25.89 22.38 27.33 1.58 7.23 16.57 34.41
TREC-COVID 64.60 62.42 70.80 3.95 29.94 52.30 70.12
SCIDOCS 14.24 10.42 11.32 0.59 6.95 11.40 17.82
SciFact 52.39 51.50 55.92 5.37 33.42 60.80 64.34
Touché-2020 26.68 27.44 26.89 1.06 6.82 33.59 25.35
NQ 39.15 44.10 44.10 0.49 9.36 16.79 53.04
FEVER 66.91 32.87 39.36 0.19 8.27 40.57 71.82
Climate-FEVER 23.71 5.93 8.22 0.28 3.72 14.00 25.66
CC News Retrieval 34.40 20.81 35.00 0.52 5.63 0.01 31.73
Sangraha-IR 44.85 41.12 47.91 3.76 21.76 30.92 42.36
MIRACL 59.34 58.11 59.24 0.69 13.76 40.98 53.56
IndicQARetrieval 69.92 67.11 67.11 21.28 46.85 74.02 61.94
mMARCO 29.49 29.94 30.92 0.48 6.98 16.53 33.03
WikiPediaRetrieval 87.38 84.40 86.71 0.03 61.28 82.54 85.65
Average 46.18 40.57 44.37 3.44 19.66 35.65 48.57
Table 2: NDCG@10 scores of the existing multilingual model and best-performing NLLB-E5 on Hindi-BEIR datasets. The best-performing model has been highlighted as bold while the second-best model has been underlined.
6.3 Baseline Models
We evaluate our Hindi-BEIR dataset on the following models:
  1. 1.
    The NLLB-E5 variations: We develop and evaluate NLLB-E5 using 600M, 1.3B and 3.3B parameter versions of the NLLB model and compare their performance [^9] https://huggingface.co/facebook/nllb-200-distilled-1.3B, https://huggingface.co/facebook/nllb-200-distilled-600M, https://huggingface.co/facebook/nllb-200-3.3B.
  2. 2.
    The BGE-M3 Chen et al. (2024) model was pre-trained on a large multilingual and cross-lingual unsupervised data and subsequently fine-tuned on high-quality multilingual retrieval datasets using a custom loss function based on the InfoNCE loss function. BGE-M3 supports a context length of 8,192.
  3. 3.
    Multilingual E5 (mE5) Wang et al. (2024c) was developed by continually pre-training the E5 model Wang et al. (2024b) on a large multilingual corpus using a weakly supervised contrastive pretraining method with InfoNCE contrastive loss van den Oord et al. (2019). It was then fine-tuned on high-quality labelled multilingual datasets for retrieval tasks. mE5 has a context length of 512 tokens.
  4. 4.
    LASER Artetxe and Schwenk (2019) focuses on universal language agnostic sentence embeddings across 93 languages. LASER uses a language-agnostic BiLSTM encoder architecture trained on parallel corpora from different languages without any retrieval-specific fine-tuning.
  5. 5.
    LaBSE Feng et al. (2022) uses a dual encoder model using BERT to obtain language-agnostic sentence embedding without any retrieval-specific fine-tuning. It has a maximum context length of 256 tokens.
  6. 6.
    BM-25 Amati (2009) is a ranking function in information retrieval based on estimating relevance by combining term frequency and document length normalization.
We extend the MTEB benchmark code repository [^10] https://github.com/embeddings-benchmark/mteb to include Hindi-BEIR and evaluate the models discussed above.
7 Results and Analysis
In this section, we compare the performance of the aforementioned baselines with our proposed NLLB-E5 model on Hindi-BEIR benchmark and highlight our key findings from the results.
Dataset Name
600M
1.3B
1.3B 3.3B
Arguana 55.23 57.24 56.50 57.68
FiQa-2018 32.83 34.19 33.92 34.41
TREC-COVID 70.07 70.53 70.54 70.12
SCIDOCS 17.53 18.23 17.67 17.82
SciFact 63.93 64.36 64.35 64.34
Touché-2020 25.89 25.21 26.13 25.35
NQ 50.14 53.38 52.77 53.04
FEVER 66.79 71.04 69.02 71.82
Climate-FEVER 25.12 22.51 23.79 25.66
CC News Retrieval 29.11 31.91 30.93 31.73
Sangraha-IR 40.51 43.30 42.84 42.36
MIRACL 50.38 52.96 52.88 53.56
IndicQARetrieval 60.47 62.10 61.69 61.94
mMARCO 31.63 34.03 33.15 33.03
WikiPediaRetrieval 84.89 86.20 86.22 85.65
Average 46.97 48.48 48.16 48.57
Table 3: NDCG@10 scores of NLLB-E5 models using NLLB encoders of different parameter sizes on Hindi-BEIR benchmark
NLLB-E5 outperforms multilingual retrievers on Hindi-BEIR: Table 2 shows NLLB1.3B-E5 model achieves better average compared to all other baselines on Hindi-BEIR. This is significant because we had state-of-the-art multilingual retriever models such as BGE-M3 and mE5 as our baselines in addition to LASER and LaBSE. It is important to note here that all these multilingual embedding models have seen a huge amount of multilingual data in their training, including Hindi. In contrast, NLLB-E5 is tuned only on English language data for learning retrieval alignment multilingual embeddings from NLLB encoder.
As we can see in Table 2 NLLB-E5 performs the best on almost all of BEIR subsets of Hindi-BEIR. More specifically, the NLLB-E5 model attains the highest scores on the Hindi versions of ArguAna, FiQA-2018, NQ, SCIDOCS and SciFact, FEVER, Climate-FEVER and mMARCO and performs 2nd best on Hindi TREC-COVID dataset. In comparison to its direct competitor, mE5-Large, the NLLB-E5 model demonstrates a substantial edge in fact-checking datasets such as FEVER, Climate-FEVER and SciFact, resulting in a significant boost of 82.47 %, 212 % and 15.06 % on the respective datasets. For non-BIER subsets too, NLLB1.3B-E5 easily outperforms LASER and LaBSE and performs competitively with mE5 and BGE-M3.
NLLB-E5 outperforms BM25 on Hindi-BEIR: BM25 is known to be a strong baseline on BEIR Thakur et al. (2021) benchmark. In fact, as seen in Table 2, BM25 indeed acts as a better baseline than LASER and LaBSE too. However, NLLB-E5 outperforms BM25 on 13/15 datasets in the Hindi-BEIR benchmark, achieving a gain of ${}+37\%$ on average. This gain justifies that we can treat NLLB-E5 as a robust neural model much better than lexical retrievers such as BM25.
BM25 has very poor performance on CC News Retrieval datasets only. The CC News Retrieval dataset was intentionally designed to test the ability of multilingual dense retrieval models to learn language-agnostic embeddings for retrieval. This dataset presents a scenario where the query is in English, and the corpus is in Hindi, which needs to move beyond token-level matching. This explains the poor performance of BM25 due to the lack of overlap between English and Hindi tokens.
Variation in Performance Across Different Tasks and Domains: One of the important aspects of building Hindi-BEIR was to see how a retriever adapts to different domains and tasks. As we can see, mE5 specifically struggles in fact-checking datasets, while BGE-M3 and NLLB-E5, although not great, still fare better for fact-checking tasks. All the models perform quite poorly for the citation prediction tasks on SCIDOCS. Also, for niche domains such as Finance and Climate, we see a sharp drop for all the models, prompting the need for focused research in these areas.
Effect of Encoder Param Size: Table 3 presents the NDCG@10 scores for the NLLB model’s encoder across different parameter sizes. As the table shows, performance generally improves as the parameter size increases, though minor anomalies exist across specific datasets. Notably, the performance gain plateaus beyond the 1.3B parameter model, with only a marginal improvement of 0.09% when comparing the NLLB-E5 model’s 1.3B (distilled) encoder to that of the 3.3B parameter version. This possibly indicates that NLLB-E5 is not lacking in model capacity yet and thus can possibly be improved further by showing more volumes of data from diverse task and domains,
8 Conclusion and Future Work
In this work, we introduced Hindi-BEIR, a comprehensive benchmark for Hindi language information retrieval. The benchmark consists of 15 datasets, with a corpus size exceeding 27 million and a query size of over 200K, covering a diverse range of 7 tasks. This provides the first-of-its-kind robust benchmark for assessing the performance of retriever models in Hindi. Our experimental results offer insights into the strengths and limitations of current multilingual retrievers on Hindi-BEIR, indicating the pressing need for further research in this domain. Additionally, we also proposed the NLLB-E5 model, a novel architecture for multilingual retriever tuning without needing multilingual training data, capable of handling data from over 200 languages, including Hindi, which outperformed existing multilingual models on Hindi-BEIR.
Future work will focus on expanding the Hindi-BEIR benchmark to include more diversity by curating additional domains such as Law and Medicine. We also plan to extend this benchmark to cover languages beyond Hindi and explore alternative multilingual encoders to optimize representation. We believe our work will have a lasting impact in developing inclusive and scalable information retrieval systems across diverse languages.
9 Limitations
While the Hindi-BEIR Benchmark and NLLB-E5 model provides valuable advancements, we acknowledge several limitations. One limitation of the Hindi-BEIR Benchmark is that it may not fully capture the breadth of tasks where retrieval models are essential. The current scope, while extensive, lacks coverage in critical domains such as Law and Medicine, which are key areas we plan to include in future expansions. Additionally, although designed primarily for Hindi, the benchmark’s extension to other languages is an area that remains unexplored in this work.
The NLLB-E5 model, while showing promise, exhibits sub-optimal performance on certain English tasks. This limitation indicates a need to explore alternative multilingual encoders to improve language performance. Additionally, we realise that the current model would architecture struggle with long texts due to a context length limitation of 508 tokens, which poses challenges in tasks requiring extended context handling. Furthermore, our approach to hyper-parameter tuning warrants further investigation to optimize model convergence and performance. Addressing these limitations will be crucial to the future development of more effective retrieval models and evaluation frameworks.
Appendix A Appendix
A.1 Dataset License
The datasets will be publicly released with the same license as the parent dataset and will be made available for research purposes.
A.2 Dataset Description
A.2.1 ArguAna
  1. 1.
    Task definition: Derived from the work by Wachsmuth et al. (2018) , the task is to retrieve the best counterargument, given an argument. Translations of arguments and counter-arguments from online debates constitute the corpus, while the translations of arguments in the original test split, after going through the filtration process based on Chrf++ scores (refer to 3.1), constitute the queries.
  2. 2.
    Domain : Misc.
An example of a query with its corresponding golden corpus has been provided in Figure 2
Figure 2: An example of a query with its corresponding golden corpus from the ArguAna Dataset
Distribution of the number of words in the corpus and queries in the ArguAna dataset has been shown in Figure 4 and Figure 3, respectively.
Figure 3: Distribution of the number of words in the queries of ArguAna Dataset
Figure 4: Distribution of Number of Words in the corpus of ArguAna Dataset
A.2.2 FiQA-2018
  1. 1.
    Task Definition: It deals with Opinion-Based Question answering. Based on the works of Maia et al. (2018),translation of the financial data extracted by crawling StackExchange posts under the Investment topic from 2009-2017, after passing through filteration processes mentioned in 3.1, acts as the corpus. While translations from the original training split acts as the queries.
  2. 2.
    Domain : Finance
An example of query with its corresponding golden corpus from the FiQA-2018 dataset has been provided in Figure 5
Figure 5: An example of a query with its corresponding golden corpus from the FiQA-2018 Dataset
Distribution of the number of words in the corpus and queries in FiQA-2018 dataset has been shown in Figure 7 and Figure 6 respectively.
Figure 6: Distribution of the number of words in the queries of FiQA-2018 Dataset
Figure 7: Distribution of Number of Words in corpus of FiQA-2018 Dataset
A.2.3 TREC-COVID
  1. 1.
    Task Definition: Voorhees et al. (2021) introduced the original TREC-COVID dataset which is an ad-hoc seach challenge based on CORD-19 dataset containing articles about the COVID-19 Pandemic. The translated and filtered version of the CORD-19 Dataset constitutes the corpus while the final cumulative judgements with query descriptions from the original task are the queries in the TREC-COVID Dataset.
  2. 2.
    Domain : Bio-Medical
An example of query with its corresponding golden corpus from the TREC-COVID dataset has been provided in Figure 8
Figure 8: An example of a query with its corresponding golden corpus from the TREC-COVID Dataset
Distribution of the number of words in the corpus and queries in TREC-COVID dataset has been shown in Figure 10 and Figure 9 respectively.
Figure 9: Distribution of the number of words in the queries of TREC-COVID Dataset
Figure 10: Distribution of Number of Words in the corpus of TREC-COVID Dataset
A.2.4 SCIDOCS
  1. 1.
    Task Definition: Inspired by Cohan et al. (2020), in this task, we expect the model to retrieve cited papers for a given scientific paper abstract as input. The corpus contains about 22k translated and filtered scientific paper abstracts and 850 translated paper titles as queries.
  2. 2.
    Domain : Scientific
An example of query with its corresponding golden corpus from the SCIDOCS dataset has been provided in Figure 11
Figure 11: An example of a query with its corresponding golden corpus from the SCIDOCS Dataset
Distribution of the number of words in the corpus and queries in the SCIDOCS dataset has been shown in Figure 13 and Figure 12, respectively.
Figure 12: Distribution of the number of words in the queries of SCIDOCS Dataset
Figure 13: Distribution of Number of Words in the corpus of SCIDOCS Dataset
A.2.5 SciFact
  1. 1.
    Task Definition: This task involves the verification of scientific claims given the abstract of scientific articles from recent literature. For this task, the model is expected to retrieve relevant abstracts with a given claim as input.
  2. 2.
    Domain : Scientific
An example of query with its corresponding golden corpus from the SciFact dataset has been provided in Figure 14
Figure 14: An example of a query with its corresponding golden corpus from the SciFact Dataset
Distribution of the number of words in the corpus and queries in the SciFact dataset has been shown in Figure 16 and Figure 15, respectively.
Figure 15: Distribution of the number of words in the queries of SciFact Dataset
Figure 16: Distribution of Number of Words in the corpus of SciFact Dataset
A.2.6 Touché-2020
  1. 1.
    Task Definition: Similar to ArguAna this task deals with this task deals with the retrieval of conversational arguments. The translated and filtered conclusion forms the title and premise for arguments Wachsmuth et al. (2017) constitutes the corpus. The translations of the Touché-2020 task data are the queries.
  2. 2.
    Domain : Miscellaneous
An example of query with its corresponding golden corpus from the Touché-2020 dataset has been provided in Figure 17
Figure 17: An example of a query with its corresponding golden corpus from the Touché-2020 Dataset
Distribution of the number of words in the corpus and queries in the Touché-2020 dataset has been shown in Figure 19 and Figure 18 respectively.
Figure 18: Distribution of the number of words in the queries of Touché-2020 Dataset
Figure 19: Distribution of Number of Words in the corpus of Touché-2020 Dataset
A.2.7 NQ
  1. 1.
    Task Definition: The task here is to retrieve the correct answer to a given question from a broad corpus of candidate answers. The NQ dataset in the Hindi-BEIR Benchmark is the translated and filtered version of the NQ dataset released by Thakur et al. (2021), which contains Google search queries and documents with paragraphs and answer spans within Wikipedia articles as the corpus.
  2. 2.
    Domain : WikiPedia
An example of a query with its corresponding golden corpus from the NQ dataset has been provided in Figure 20
Figure 20: An example of a query with its corresponding golden corpus from the NQ Dataset
Distribution of the number of words in the corpus and queries in the NQ dataset has been shown in Figure 22 and Figure 21, respectively.
Figure 21: Distribution of the number of words in the queries of NQ Dataset
Figure 22: Distribution of Number of Words in the corpus of NQ Dataset
A.2.8 FEVER
  1. 1.
    Task Definition: Similar to SciFact, here, the task is to retrieve relevant documents that claim a given fact (query). We translate and filter the test split of the FEVER dataset as proposed by Thakur et al. (2021) and include it in the Hindi-BEIR Benchmark.
  2. 2.
    Domain : Wikipedia
An example of a query with its corresponding golden corpus from the FEVER dataset has been provided in Figure 23
Figure 23: An example of a query with its corresponding golden corpus from the FEVER Dataset
Distribution of the number of words in the corpus and queries in the FEVER dataset has been shown in Figure 25 and Figure 24, respectively.
Figure 24: Distribution of the number of words in the queries of FEVER Dataset
Figure 25: Distribution of Number of Words in the corpus of FEVER Dataset
A.2.9 Climate-FEVER
  1. 1.
    Task Definition: Similar to FEVER, Climate-FEVER is a dataset for the verification of real-world climate claims. We translate and filter the test split of the Climate-FEVER dataset as proposed by Thakur et al. (2021) and include it in the Hindi-BEIR Benchmark.
  2. 2.
    Domain : Wikipedia
An example of query with its corresponding golden corpus from the Climate-FEVER dataset has been provided in Figure 26
Figure 26: An example of a query with its corresponding golden corpus from the Climate-FEVER Dataset
Distribution of the number of words in the corpus and queries in the Climate-FEVER dataset has been shown in Figure 28 and Figure 27, respectively.
Figure 27: Distribution of the number of words in the queries of Climate-FEVER Dataset
Figure 28: Distribution of Number of Words in corpus of Climate-FEVER Dataset
A.2.10 CC News Retrieval
  1. 1.
    Task Definition: CC News Retrieval Introduces the task to retrieving relevant news articles given a news title. Subsection 3.2 talks in detail about the data creation process.
  2. 2.
    Domain : News
An example of a query with its corresponding golden corpus from the CC News Retrieval dataset has been provided in Figure 29
Figure 29: An example of a query with its corresponding golden corpus from the CC News Retrieval Dataset
Distribution of the number of words in the corpus and queries in CC News Retrieval dataset has been shown in Figure 31 and Figure 30 respectively.
Figure 30: Distribution of the number of words in the queries of CC News Retrieval Dataset
Figure 31: Distribution of Number of Words in corpus of CC News Retrieval Dataset
A.2.11 Sangraha-IR
  1. 1.
    Task Definition: Sangraha-IR introduces the task to retrieving relevant text documents given a question in Hindi. Subsection 3.2 talks in detail about the data creation process.
  2. 2.
    Domain : Miscellaneous
An example of a query with its corresponding golden corpus from the Sangraha-IR dataset has been provided in Figure 32
Figure 32: An example of a query with its corresponding golden corpus from the Sangraha-IR Dataset
Distribution of the number of words in the corpus and queries in the Sangraha-IR dataset has been shown in Figure 34 and Figure 33, respectively.
Figure 33: Distribution of the number of words in the queries of Sangraha-IR Dataset
Figure 34: Distribution of Number of Words in corpus of Sangraha-IR Dataset
A.2.12 MIRACL
  1. 1.
    Task Definition: MIRACL has been created from the Wikipedia dump of each language. Only plain text is considered, while images, etc., are omitted. The plain text is split up into multiple paragraphs, which act as the corpus. We consider the dev split of the Hindi subset of the original MIRACL in the Hindi-BEIR Benchmark.
  2. 2.
    Domain : Wikipedia
An example of query with its corresponding golden corpus from the MIRACL dataset has been provided in Figure 35
Figure 35: An example of a query with its corresponding golden corpus from the MIRACL Dataset
Distribution of the number of words in the corpus and queries in the MIRACL dataset has been shown in Figure 37 and Figure 36 respectively.
Figure 36: Distribution of the number of words in the queries of MIRACL Dataset
Figure 37: Distribution of Number of Words in the corpus of MIRACL Dataset
A.2.13 IndicQARetrieval
  1. 1.
    Task Definition: It is created by transforming the IndicQA dataset  [^11] https://huggingface.co/datasets/ai4bharat/IndicQA to a retrieval dataset. The task is similar to NQ, where the model is expected to return the paragraph that contains the answer to the question when given a question.
  2. 2.
    Domain : Wikipedia
An example of query with its corresponding golden corpus from the IndicQARetrieval dataset has been provided in Figure 38
Figure 38: An example of a query with its corresponding golden corpus from the IndicQARetrieval Dataset
Distribution of the number of words in the corpus and queries in the IndicQARetrieval dataset has been shown in Figure 40 and Figure 39 respectively.
Figure 39: Distribution of the number of words in the queries of IndicQARetrieval Dataset
Figure 40: Distribution of Number of Words in the corpus of IndicQARetrieval Dataset
A.2.14 mMARCO
  1. 1.
    Task Definition: It is a multilingual version of the MSMARCO dataset. The dataset contains the translation of queries from Bing search logs with one text passage from various web sources annotated as relevant.
  2. 2.
    Domain : Miscellaneous
An example of a query with its corresponding golden corpus from the mMARCO dataset has been provided in Figure 41
Figure 41: An example of a query with its corresponding golden corpus from the mMARCO Dataset
Distribution of the number of words in the corpus and queries in the mMARCO dataset has been shown in Figure 43 and Figure 42 respectively.
Figure 42: Distribution of the number of words in the queries of mMARCO Dataset
Figure 43: Distribution of Number of Words in the corpus of mMARCO Dataset
A.2.15 WikiPediaRetreival
  1. 1.
    Task Definition: It is similar to a question-answering task dataset like NQ, where given a query, the model is expected to retrieve a relevant article which answers the question. We have included the Hindi subset of WikiPediaRetrieval dataset  [^12] https://huggingface.co/collections/ellamind/mmteb-6661723dc229e1da8e837cdf
  2. 2.
    Domain : Wikipedia
An example of a query with its corresponding golden corpus from the WikiPediaRetreival dataset has been provided in Figure 44
Figure 44: An example of a query with its corresponding golden corpus from the WikiPediaRetreival Dataset
Distribution of the number of words in the corpus and queries in the WikiPediaRetreival dataset has been shown in Figure 46 and Figure 45 respectively.
Figure 45: Distribution of the number of words in the queries of WikiPediaRetreival Dataset
Figure 46: Distribution of Number of Words in the corpus of WikiPediaRetreival Dataset
A.3 Why choose IndicTrans2 over other available translation models?
Ans: Gala et al. (2023) clearly illustrate the superior performance of IndicTrans2 over other models and systems like NLLB and Google Translate for English to Hindi tasks. In our preliminary analysis on a subset of the BEIR datasets (Chrf scores show in Table 4 ), we also observed that IndicTrans2 outperformed alternative models, such as NLLB, in terms of translation quality.
Dataset IndicTrans2 NLLB
Arguana 56.30 37.12
NQ 76.13 60.92
Table 4: Chrf scores between the original English text and back translated english text by the respective model for 20,000 randomly selected samples from each dataset.
A.4 Prompts used for Sangraha-IR Query Generation
The prompt that was used to obtain answers from the Gemini-Flash-1.5 Models is as follows
Given the above text, generate a question in Hindi which can only be answered by the above passage. The question should be difficult with only semantic similarity and should not contain any lexical overlap; that is, do not directly use the exact phrases as shown in the above passage. You can use synonyms and make the question as complex as possible. Only return the relevant question and nothing else in the format given below : $<$QUESTION$>$: …question in hindi…. </absent</< /QUESTION$>$
A.5 Ablation of LoRA parameters
Dataset Name
NLLB1.3B(dist)-E5
(W/O LoRA, ST)
NLLB1.3B(dist)-E5
(W LoRA, ST)
ArguAna 57.24 58.19
FiQA-2018 34.19 34.46
TREC-COVID 70.53 70.52
SCIDOCS 18.23 18.06
SciFact 64.36 65.23
Touché-2020 25.21 25.74
NQ 53.38 51.09
FEVER 71.04 72.71
Climate-FEVER 22.51 23.63
CC News Retrieval 31.91 29.47
MIRACL 52.96 49.76
IndicQARetrieval 62.10 61.81
mMARCO 34.03 31.78
WikiPediaRetrieval 86.20 84.36
Average 48.84 48.34
Table 5: NDCG@10 scores of NLLB-E5 model with and without LoRA adapters on E5 model on the Hindi-BEIR Benchmark.
Similar to Schmidt et al. (2024), we also experiment with the addition of LoRA adapters during training on top of the E5 model. The LoRA adapters were applied to the "query", "key", and "value" layers of the retrieval model head, with a rank of 32, an alpha parameter of 64, and a dropout rate of 0.05.
Table 5 presents a side-by-side comparison of the NLLB-E5 model with and without LoRA adapters applied to the E5 head. The results indicate a slight performance boost on certain datasets, particularly those derived from the original BEIR benchmark, which closely resembles the training data. However, it is important to note a significant drop in performance on non-BEIR datasets, such as CC-News Retrieval, MIRACL, IndicQA Retrieval, mMARCO, and Wikipedia Retrieval. We hypothesize that this drop could be due to the LoRA adapters causing the model to overfit on datasets in the style of Sentence Transformers, thereby reducing its generalization capabilities. The overall stronger performance of the model without LoRA adapters is further supported by its higher average NDCG@10 score, and thus, we decided to proceed with the "without LoRA" model setup.